OpenStudy (anonymous):

Find points on the curve x^2 - xy + y^2 = 9 that are closest to and furthest from the origin.

OpenStudy (anonymous):

Go to polar coordinates!

ganeshie8 (ganeshie8):

you familiar wid lagrange ?

ganeshie8 (ganeshie8):

okay distance function is simply : x^2+y^2

OpenStudy (anonymous):

Going that way, you need to minimize $x^2+y^2$ subject to the constraint that $x^2-xy+y^2-9=0$

OpenStudy (anonymous):

I have differentiated implicitly and gy/dx= 2x - y / x - 2y. The next step is to write down a function that represents the distance of a point from the origin. I am stuck.

OpenStudy (anonymous):

define your function to be the distance function and subtract the constraint multiplied by the lagrangian multiplier. then diffeentiate with respect to x, y and the multiplier.

OpenStudy (anonymous):

.....easier, though is to go to polar coordinates. Then you can avoid constraints altogether.

OpenStudy (anonymous):

we have not yet covered the above methods so there must be a simpler way. step 1 is to implicitly differentiate which I have done step 2 is to write down a function that represents the distance of a point from the origin and to differentiate D(x) step 3 is to solve dy/dx=0, d'(x)=0 and x^2 - xy + y^2 = 9 together to determine the required points of the curve. I am stuck on step 2

OpenStudy (anonymous):

What are you differentiating and with respect to which variable?

OpenStudy (anonymous):

you are differentiating the distance function. I am not sure what the distance function is.

OpenStudy (anonymous):

So you need to start with step 2

ganeshie8 (ganeshie8):

distance between (0, 0 ) and any point on curve (x,y ) is : sqrt(x^2+y^2) you may as well optimize its square : x^2+y^2

OpenStudy (anonymous):

ok thanks I will try work it out and come back with my solution to see if im on the right track

ganeshie8 (ganeshie8):

still stuck on this ?

OpenStudy (anonymous):

sqrt(x^2+y^2) and x^2+y^2 has extrema for the same values of x and y. The latter is easier to differentiate with than the former. So in other words, we need to find extrema of x^2+y^2. One approach is to solve the constraint for y and eliminate y from this expression. Then you get an expression that only has x in it, and you can differentiate to find minima. Easier is to substitute x=r cos(theta) and y=r sin(theta) into the equation for the curve. This results in a very simple equation for r^2 that you can easily minimize.

OpenStudy (anonymous):

Ok I am lost. I am using the steps provided see above. So step 2 I have the distance function and to take it's derivative. If I use the method of taking it's square I get d'=-x/y So now I am up to step 3. Making sense?

OpenStudy (anonymous):

There are several approaches: (1) in the equation for the curve, substitute x = r cos(theta), y=r sin(theta) and solve for r^2 and then look for max/min of r^2 as a function of theta. This is the easiest approach. (2) Solve the equation for the curve for y in terms of x and substitute this into x^2 + y^2. This will give you an expression for the distance function in terms of x only. Differentiate it to find its max/min. (3) use lagragian multipliers. (This is out since you do't know this method.)

OpenStudy (ranga):

x^2 - xy + y^2 = 9 ----- (1) To find the maximum/minimum distance of a point (x,y) on the curve from the origin we find the max/min of x^2 + y^2 (distance squared). Let D = x^2 + y^2 Differentiate with respect to x: D' = 2x + 2yy' = 0 2x = -2yy' x = - yy' ----- (2) Find y' by implictly differentiating the original equation: x^2 - xy + y^2 = 9 2x - xy' - y + 2yy' = 0 2x - y + y'(2y - x) = 0 y' = -(2x-y)/(2y-x) ---- (3) Put y' from (3) into (2): x = -y * -(2x-y)/(2y-x) = y(2x-y)/(2y-x) x(2y-x) = y(2x-y) 2xy - x^2 = 2xy - y^2 x^2 = y^2 y = x or -x Put y = x in (1): x^2 - x^2 + x^2 = 9 x^2 = 9 x = -3 or +3 y = -3 or +3 The two points are: (-3,3) and (3,3) Put y = -x in (1): x^2 + x^2 + x^2 = 9 3x^2 = 9 x^2 = 3 x = -sqrt(3) or sqrt(3) y = +sqrt(3) oe -sqrt(3) The two points are: (-sqrt(3),sqrt(3)) and (sqrt(3),-sqrt(3)) Substitute the points (-3,3) and (3,3) into the distance-squared equation: x^2 + y^2 = 9 + 9 = 18 distance = sqrt(18) = sqrt(2*9) = 3*sqrt(2) Both points give the same distance of 3*sqrt(2) Substitute the points (-sqrt(3),sqrt(3)) and (sqrt(3),-sqrt(3))into the distance-squared equation: x^2 + y^2 = 3 + 3 = 6 distance = sqrt(6) Both points give the same distance of sqrt(6) Closest points are: (-sqrt(3),sqrt(3)) and (sqrt(3),-sqrt(3)) Farthest points are: (-3,3) and (3,3) Minimum distance = sqrt(6) Maximum distance = 3*sqrt(2)

OpenStudy (anonymous):

The solution in polar coordinates: let x =r cos(t), y = r sin(t). Plugging into the curve: x^2 - xy + y^2 = 9 we get r^2 - r^2 sin(t) cos(t) = 9 r^2 (1 - 0.5 sin(2t)) = 9 r^2=9/(1 - 0.5 sin(2t)) = Hence r^2 has max when sin(2t) = 1 i.e. when t = pi/4 or 5pi/4 that is r^2 = 18 and x = y = 3 or x = y = -3. r^2 has min when sin(2t) = -1 i.e. when t = -pi/4 or -5pi/4 that is r^2 = 6 and x = -y = sqrt(3) or x = -y = -sqrt(3).

OpenStudy (anonymous):

Thank you so much for your help I looked for similar examples but it was hard to find.

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

Thanks ljensen for the polar coordinates example. A lot simpler but we have not yet covered that topic so I can't use it :(